# Adaptive Local Neighborhood-based Neural Networks for MR Image Reconstruction from Undersampled Data

### v2

@@ I Introduction
In applications like X-ray computed tomography (CT) [1] and magnetic resonance imaging (MRI) [2], reconstructing images from undersampled or corrupted observations is of critical importance. For example, this is necessary to reduce a patient’s exposure to radiation in CT or reduce time spent acquiring MRI data. MRI scans involve sequential data acquisition resulting in long acquisition times that are not only a burden for patients and hospitals, but also make MRI susceptible to motion artifacts. Reconstructing images from limited measurements can speed up the MRI scan, but usually entails solving an ill-posed inverse problem. Recent approaches to accelerating MRI acquisition such as compressed sensing (CS) [3] reduce scan time by collecting fewer measurements while preserving image quality by exploiting image priors or regularizers. Historically, regularization in CS-MRI has been based on sparsity of wavelet coefficients [4] or using total variation [5]. While conventional CS assumes sparse or incoherent signals, approaches based on learned image models have been shown to be more effective for MRI reconstruction, starting with learned synthesis dictionaries [6, 7]. The dictionary parameters could be learned from unpaired clean image patches from a dataset and used for reconstruction or learned simultaneously with image reconstruction [8, 9, 10, 6]. Additionally, recent advances in sparsifying transform learning have resulted in efficient or inexpensive data-adaptive sparsity-based reconstruction frameworks for MRI [11, 12, 13]. Other contemporary techniques could allow learning explicit regularizers in a supervised manner [14] for improved image restoration.
@@
Deep learning (DL) has emerged as a potent methodology for tackling large-scale inverse problems, notably in enhancing image reconstruction techniques in MRI and CT. Predominantly, end-to-end CNN, as exemplified by the U-net model [15, 16], have been employed to mitigate artifacts arising from undersampling in MRI datasets. Additionally, a plethora of alternative network models such as the Transformer [17], and Generative Adversarial Networks (GANs)[18], have demonstrated their effectiveness in MRI reconstruction, as detailed in comprehensive reviews like [12]. Furthermore, transfer learning [19] has also been used with neural networks for MRI reconstruction to achieve domain transfer.
@@
To enhance both stability and performance, hybrid-domain approaches such as [20] enforce data consistency (i.e., the reconstruction is enforced to be consistent with the measurement model) all through training and reconstruction. Networks incorporating data consistency layers are pivotal in MR imaging, maintaining alignment between the reconstructed image and the original data in k-space [21, 22]. This category encompasses various methodologies, including deep unrolling-based methods [23, 24](which adapt traditional iterative algorithms to learn regularization parameters) regularization by denoising approaches [25], and plug-and-play methods [26], among others. Distinctively, the ADMM-CSNet [23] utilizes neural networks for the optimization of ADMM parameters, diverging from the ISTA-Net [27], which focuses on refining CS reconstruction models grounded in the Iterative Shrinkage-Thresholding Algorithm. While these CNN-based reconstruction methods have demonstrated superiority over traditional CS techniques, concerns regarding their stability and interpretability persist, as highlighted in [28].
@@
Apart from algorithmic advances, another driving force behind deep learning-based reconstruction is the the rapid growth of publicly available training datasets. The availability of (paired or unpaired) training data sets made possible by efforts like OCMR [29] and fastMRI [30] has enabled rapidly demonstrating the capacity of deep learning-based algorithms for improved image reconstruction or denoising quality in MRI applications.
@@
However, one major drawback of these learned approaches is that they typically require large training datasets such as fully sampled MRI data to be effective. A recent scan-specific deep learning method is the deep image prior [31], which has been applied to MRI [32] and learns a neural network for reconstruction in an unsupervised fashion from a single image’s measurements. Other scan-adaptive methods include RAKI [33], which is a nonlinear deep learning-based auto-regressive auto-calibrated reconstruction method. RAKI could be viewed as a deep neural network-based version of the parallel imaging scheme GRAPPA [34]. LORAKI [35] is another scheme that trains an autocalibrated recurrent neural network (RNN) to recover missing k-space data. All these methods learn scan-specific networks without requiring large datasets. A related approach dubbed self-supervised learning has also shown promise for MRI [36] and uses a large unpaired data set.
@@ I-A Contributions
  • The proposed models are trained using a small number of adaptively chosen neighbors that are in proximity (or are similar to in a sense) to the underlying (to be reconstructed) image (cf. [37] for a slightly related approach in the context of patch-based dictionary learning).
@@
  • Our experimental results on multiple datasets (fastMRI, Stanford FSE, and fastMRI+) and across multiple k-space undersampling factors showed that the proposed local adaptation techniques surpass networks trained globally on larger datasets. We demonstrated improved performance against scan-specific deep learning methods such as deep image prior, RAKI, and LORAKI, even when using a small number of neighbors for training.
@@ II-A Multi-coil MRI Reconstruction
When an image $\bm{x}\in\mathbb{C}^{q}$ (vectorized) is sufficiently sparse in some transform domain and the transform is sufficiently incoherent with the measurement operator, the theory of compressed sensing [3, 38] enables accurate image recovery from limited measurements. The image reconstruction problem in MRI is typically formulated as an optimization of a data-fidelity penalty and a regularizer as follows:
@@ II-B Neighbor Search
As a result, we used different metrics for evaluating our approach in this work. The distances serve as a proxy for data similarity, with nearby data considered similar and distant data considered dissimilar. We used the Euclidean distance, Manhattan distance, and normalized cross-correlation as distance metrics as follows.
@@
$d^{NCC}(\bm{x},\bm{x}_{n})=\frac{\begin{vmatrix}\bm{x}^{H}\bm{x}_{n}\end{vmatrix}}{\|\bm{x}\|_{2}\,\|\bm{x}_{n}\|_{2}}$
In all cases, we select the top $k$ most similar neighbors from a set that correspond to the $k$ smallest distances in (2). The indices of the chosen images are in the set $\hat{C}_{\bm{x}}$, i.e., they are the minimizer in (2). These neighbors can be used to train the local model. These are expected to capture structures most similar to the image being reconstructed, enabling a highly effective reconstruction model to be learned.
@@ IV-A Experimental Setup
Datasets & Models: We evaluated the effectiveness of the proposed LONDN-MRI reconstruction method on multiple datasets: the multi-coil fastMRI knee and brain datasets [30, 40], the fastMRI+ dataset [^1] https://github.com/microsoft/fastmri-plus/tree/main (which is just an annotated version of fastMRI indicating pathologies), and the Stanford 2D FSE [41] dataset. The results obtained on the fastMRI knee dataset and the Stanford FSE data are described in Section IV-B. The fastMRI brain and fastMRI+ data are used in the studies in Section IV-D. For training, we randomly selected a subset of $3000$ images from the fastMRI knee and brain datasets and the same for the fastMRI+ case. We used $2000$ training images for the smaller Stanford FSE dataset. We used $15$ or $20$ images for testing in different scenarios, which were randomly chosen. In some experiments, we evaluated the effect of training set size, where we worked with fewer or more images in the training set. Coil sensitivity maps for model-based reconstruction were generated for each scan using the BART toolbox [42]. We tested obtaining these using either the fully-sampled k-space or only center of k-space data and noticed very little difference in reconstruction quality between the two approaches.
@@
Since the proposed LONDN-MRI framework is quite general and can be combined with any supervised deep learning based reconstruction approach, we chose the recent popular model-based deep learning (MoDL) reconstruction network and compared globally (over large set of training samples) and locally (over very small matched set of samples) learned versions of the model for different choices of deep denoisers in the network.222 See https://github.com/sjames40/Multi_coil_local_model for our code in PyTorch. We performed reconstructions at fourfold or 4x acceleration (25.0% sampling) as well as at eightfold or 8x acceleration (12.5% sampling) of the k-space acquisition. In all cases, variable density 1D random Cartesian (phase-encode) undersampling of k-space was performed. The initial image estimates for MoDL were obtained by applying the adjoint of the measurement operator to the subsampled k-space data, and were then used to train both local and global versions of MoDL networks. In our local versions (LONDN-MRI), we used 30 images for training (searched from e.g., 3000 images). while the global versions used the full subset of training images.
@@
Network Architectures & Training: We trained two types of MoDL models at 4x and 8x k-space undersampling, respectively. One used the well-known UNet denoiser, with a two-channel input and two-channel output, where the real and imaginary parts of an image are separated into two channels. The network weights during training were initialized randomly (normally distributed). The ADAM optimizer was utilized for training the network weights. For LONDN-MRI, we used an initial learning rate of $6\times 10^{-5}$ with a multi-step learning rate scheduler, which decreases the learning rate at $100$ and $150$ epochs with learning rate decay $0.65$. For training globally, we used an initial learning rate of $1\times 10^{-4}$ with $150$ epochs of training and a multi-step learning rate scheduler that decreased the learning rate at $50$ and $100$ epochs with learning rate decay $0.6$. For LONDN-MRI, MoDL with $5$ iterations was used with a shallow UNet that had $2$ layers in the encoder and decoder, respectively. We used a shallow network with dropout for the local model to avoid over-fitting to the very small training set. For the MoDL network trained globally (on large dataset) for making comparisons with, we utilized $4$ layers in the decoder and encoder in UNet and $6$ MoDL blocks. We used a batch size of $2$ during training for both the global and local cases. Furthermore, for the data-consistency term, we used a tolerance of $10^{-5}$ in CG and a $\mu/\nu$ ratio of 0.1. Also, we chose the regularization weight $\lambda$ as $10^{-9}$ for LONDN-MRI, unless specified otherwise.
@@
For the second MoDL architecture, we used the recent state-of-the-art denoising network DIDN [43, 44]. Due to the high complexity of the DIDN network, we first pre-trained it on the larger (global) dataset (learning rate, etc., similar to the UNet case) before adapting the weights within LONDN-MRI for each scan. This is an alternative to constructing shallower versions of a network for local adaptation. The ADAM optimizer was utilized for training, with a learning rate of $5\times 10^{-5}$ in LONDN-MRI. We used $6$ iterations of MoDL with the DIDN denoiser for which we used $3$ down-up blocks (DUBs). The number of epochs for training was $30$ in LONDN-MRI. The remaining training parameters were chosen similarly as in the previous UNet-based case. Using a pre-trained state-of-the-art denoiser allows the local adaptation to converge faster.
@@ IV-B Results and Comparisons
When varying the size of the training set, the global approach was trained on the full set each time, whereas the local approach performed training on small subsets of $30$ training pairs selected from the larger datasets. The iterations of the LONDN-MRI scheme quickly improve reconstruction performance, and even with only 2 LONDN-MRI alternations, the PSNR values begin approaching the oracle setting.The LONDN schemes (oracle or iterative) consistently outperform the globally trained networks across the different training set sizes considered. We note that the results for the globally trained model with many ($6000$) training scans match closely the LONDN-MRI results, when LONDN-MRI uses a smaller overall training set ($3000$ scans) for neighbor search. This illustrates the potential of our approach with limited training data, when compared with models trained on larger sets. Figure 2 compares the SSIM and HFEN reconstruction metrics using bar graphs, where a similar trend is observed as with PSNR.
Figs. 4 and 5 show images reconstructed by different methods at 8x and 4x undersampling, respectively. The LONDN-MRI reconstructions (either iterative or oracle) show fewer artifacts, sharper features, and fewer errors than the global MoDL and initial aliased reconstructions. The iterative LONDN-MRI results are also quite close to the oracle result.
@@ IV-C Behavior of LONDN-MRI
$\text{NMA}:=\frac{100}{|C|}\sum_{r\in\mathcal{T}}\frac{|\hat{C}_{r}\cap C_{r}^{*}|}{k},$ (8)
@@
Convergence of Loss in Bilevel Optimization: Next, we study the behavior of the alternating LONDN-MRI algorithm as a heuristic for the bilevel optimization formulation in (7). Here, we used an MoDL network with the UNet denoiser and $k=30$ training pairs were chosen (from 3000 cases) in the local dataset in each iteration of LONDN-MRI. The UNet weights were randomly initialized to begin with, and the neighbor search in the first iteration of LONDN-MRI was performed using $\bm{x}^{0}$ and correspondingly generated aliased training images. Fig. 11 plots the upper-level loss in (7) (in a root mean squared error form) after each iteration of LONDN-MRI for a test image. Here, we ran many iterations to verify convergence. We observe that the loss changes very little after a few iterations and stabilizes. This matches with the behavior of the neighbor search accuracy bar plots. The result indicates that the proposed alternating scheme could be a reasonable heuristic for reducing the loss in the challenging problem (7). Finally, we compare the loss values in Fig. 11 with an oracle loss, where the upper-level loss in (7) is computed using the ground truth test image and its $k$ nearest neighbors. It is clear that the loss values in LONDN-MRI converge very close to the oracle loss, indicating the potential for our scheme.
@@
Time Consumption Trade-offs: To further understand the time efficiency of our method across different neighborhood sizes for practical applicability, we conducted comparative analyses using three models: an image-domain UNet denoiser, MODL with UNet denoiser, and the MODL with DIDN denoiser. The experiments were run on an NVIDIA GeForce RTX A5000 GPU. The PSNR vs. runtime trade-offs depicted in Figure 13 shed light on the time consumption for each model configuration. It is observed that some decrease in the number of neighbors leads to reduced time consumption without significantly compromising image quality. In addition, the results show the effectiveness of starting with a pre-trained DIDN model to improved the reconstruction, as it enhances the efficiency of the reconstruction process, reducing it to order of seconds.

### v3

@@ I Introduction
In applications like X-ray computed tomography (CT) [1] and magnetic resonance imaging (MRI) [2], reconstructing images from undersampled or corrupted observations is of critical importance. For example, this is necessary to reduce a patient’s exposure to radiation in CT or reduce time spent acquiring MRI data. MRI scans involve sequential data acquisition resulting in long acquisition times that are not only a burden for patients and hospitals, but also make MRI susceptible to motion artifacts. Reconstructing images from limited measurements can speed up the MRI scan, but usually entails solving an ill-posed inverse problem. Recent approaches to accelerating MRI acquisition such as compressed sensing (CS) [3] reduce scan time by collecting fewer measurements while preserving image quality by exploiting image priors or regularizers. Historically, regularization in CS-MRI has been based on sparsity of wavelet coefficients [4] or using total variation [5]. While conventional CS assumes sparsity of signals in known bases, approaches based on learned image models have been shown to be more effective for MRI reconstruction, starting with learned synthesis dictionaries [6, 7]. The dictionary parameters could be learned from unpaired clean image patches from a dataset and used for reconstruction or learned simultaneously with image reconstruction [8, 9, 10, 6]. Additionally, recent advances in sparsifying transform learning have resulted in efficient or inexpensive data-adaptive sparsity-based reconstruction frameworks for MRI [11, 12, 13]. Other contemporary techniques could allow learning explicit regularizers in a supervised manner [14] for improved image restoration.
@@
Deep learning (DL) has emerged as a potent methodology for tackling large-scale inverse problems, notably in enhancing image reconstruction techniques in MRI and CT. Predominantly, end-to-end CNN, as exemplified by the U-net model [15, 16], have been employed to mitigate artifacts arising from undersampling in MRI data. Additionally, a plethora of alternative network models such as the Transformer [17], and Generative Adversarial Networks (GANs) [18], have demonstrated their effectiveness in MRI reconstruction, as detailed in comprehensive reviews like [12]. Furthermore, transfer learning [19] has also been used with neural networks for MRI reconstruction to achieve domain transfer.
@@
To enhance both stability and performance, hybrid-domain approaches such as [20] enforce data consistency (i.e., the reconstruction is enforced to be consistent with the measurement model) all through training and reconstruction. Networks incorporating data consistency layers are pivotal in MR imaging for maintaining alignment between the reconstructed image and the original data in k-space [21, 22]. This category encompasses various methodologies, including deep unrolling-based methods [23, 24](which adapt traditional iterative algorithms to learn regularization parameters), regularization by denoising approaches [25], and plug-and-play methods [26], among others. Distinctively, the ADMM-CSNet [23] utilizes neural networks for the optimization of ADMM parameters, diverging from the ISTA-Net [27], which focuses on refining CS reconstruction models grounded in the Iterative Shrinkage-Thresholding Algorithm. The 1D deep low-rank and sparse network (ODLS) [28] demonstrates enhanced robustness for 2D MR image reconstruction, particularly in scenarios characterized by a limited number of training samples. While these deep learning-based reconstruction methods have demonstrated superiority over traditional CS techniques, concerns regarding their stability and interpretability persist, as highlighted in [29].
@@
Apart from algorithmic advances, another driving force behind deep learning-based reconstruction is the the rapid growth of publicly available training datasets. The availability of (paired or unpaired) training data sets made possible by efforts like OCMR [30] and fastMRI [31] has enabled rapidly demonstrating the capacity of deep learning-based algorithms for improved image reconstruction or denoising in MRI applications.
@@
However, one major drawback of these learned approaches is that they typically require large training datasets such as fully sampled MRI data to be effective. A recent scan-specific deep learning method is the deep image prior [32], which has been applied to MRI [33] and learns a neural network for reconstruction in an unsupervised fashion from a single image’s measurements. Other scan-adaptive methods include RAKI [34], which is a nonlinear deep learning-based auto-regressive auto-calibrated reconstruction method. RAKI could be viewed as a deep neural network-based version of the parallel imaging scheme GRAPPA [35]. LORAKI [36, 37] is another scheme that trains an autocalibrated recurrent neural network (RNN) to recover missing k-space data. All these methods learn scan-specific networks without requiring large datasets. A related approach dubbed self-supervised learning has also shown promise for MRI [38] and uses a large unpaired data set.
@@ I-A Contributions
  • The proposed models are trained using a small number of adaptively chosen neighbors that are in proximity (or are similar in a sense) to the underlying (to be reconstructed) image (cf. [39] for a slightly related approach in the context of patch-based dictionary learning).
@@
  • Our experimental results on multiple datasets (fastMRI, Stanford FSE, and fastMRI+) and across multiple k-space undersampling factors show that the proposed local adaptation techniques surpass networks trained globally on larger datasets. We demonstrate improved performance against scan-specific deep learning methods such as deep image prior, RAKI, and LORAKI, even when using a small number of neighbors for training.
@@ II-A Multi-coil MRI Reconstruction
When an image $\bm{x}\in\mathbb{C}^{q}$ (vectorized) is sufficiently sparse in some transform domain and the transform is sufficiently incoherent with the measurement operator, compressed sensing theory [3, 40] enables accurate image recovery from limited measurements. The image reconstruction problem in MRI is typically formulated as an optimization of a data-fidelity penalty and a regularizer as follows:
@@ II-B Neighbor Search
As a result, we used different metrics for evaluating our approach in this work. The distances serve as a proxy for data similarity, with nearby data considered similar and distant data considered dissimilar. We used the Manhattan distance, Euclidean distance, and normalized cross-correlation as distance metrics as follows.
@@
$d^{NCC}(\bm{x},\bm{x}_{n})=1-\frac{\begin{vmatrix}\bm{x}^{H}\bm{x}_{n}\end{vmatrix}}{\|\bm{x}\|_{2}\,\|\bm{x}_{n}\|_{2}}.$
In all cases, we select the top $k$ most similar neighbors (from a set) that correspond to the $k$ smallest distances in (2). The indices of the chosen images are in the set $\hat{C}_{\bm{x}}$, i.e., they are the minimizer in (2). These neighbors can be used to train the local model. These are expected to capture structures most similar to the image being reconstructed, enabling a highly effective reconstruction model to be learned.
@@ IV-A Experimental Setup
Datasets & Models: We evaluated the effectiveness of the proposed LONDN-MRI reconstruction method on multiple datasets: the multi-coil fastMRI knee and brain datasets [31, 42], the fastMRI+ dataset [^1] https://github.com/microsoft/fastmri-plus/tree/main (which is just an annotated version of fastMRI indicating pathologies), and the Stanford 2D FSE [43] dataset. The results obtained on the fastMRI knee dataset and the Stanford FSE data are described in Section IV-B. The fastMRI brain and fastMRI+ data are used in the studies in Section IV-D. For training, we randomly selected a subset of $3000$ images from the fastMRI knee and brain datasets and the same for the fastMRI+ case. We used $2000$ training images for the smaller Stanford FSE dataset. We used $15$ or $20$ images for testing in different scenarios, which were randomly chosen. In some experiments, we evaluated the effect of training set size, where we worked with fewer or more images in the training set.
Coil sensitivity maps for model-based reconstruction were generated for each scan using the BART toolbox [44]. We tested obtaining these using either the fully-sampled k-space or only center of k-space data and noticed very little difference in reconstruction quality between the two approaches.
@@
Since the proposed LONDN-MRI framework is quite general and can be combined with any supervised deep learning based reconstruction approach, we chose the recent popular model-based deep learning (MoDL) reconstruction network and compared globally (over large set of training samples) and locally (over very small matched set of samples) learned versions of the model for different choices of deep denoisers in the network 222 See https://github.com/sjames40/Multi_coil_local_model for our code in PyTorch. . We performed reconstructions at fourfold or 4x acceleration (25.0% sampling) as well as at eightfold or 8x acceleration (12.5% sampling) of the k-space acquisition. In all cases, variable density 1D random Cartesian (phase-encode) undersampling of k-space was performed. The initial image estimates for MoDL were obtained by applying the adjoint of the measurement operator to the subsampled k-space data, and were then used to train both local and global versions of MoDL networks. In our local versions (LONDN-MRI), we used 30 images for training (searched from e.g., 3000 images). while the global versions used the full subset of training images.
@@
Network Architectures & Training: We trained two types of MoDL models at 4x and 8x k-space undersampling, respectively. One used the well-known UNet denoiser, with a two-channel input and two-channel output, where the real and imaginary parts of an image are separated into two channels. The network weights during training were initialized randomly (normally distributed). The ADAM optimizer was utilized for training the network weights. For LONDN-MRI, we used an initial learning rate of $6\times 10^{-5}$ with a multi-step learning rate scheduler, which decreases the learning rate at $100$ and $150$ epochs with learning rate decay $0.65$. For training globally, we used an initial learning rate of $1\times 10^{-4}$ with $150$ epochs of training and a multi-step learning rate scheduler that decreased the learning rate at $50$ and $100$ epochs with learning rate decay $0.6$. For LONDN-MRI, MoDL with $5$ iterations was used with a shallow UNet that had $2$ layers in the encoder and decoder, respectively. We used a shallow network with dropout for the local model to avoid over-fitting to the very small training set.
For the MoDL network trained globally (on large dataset) for making comparisons with, we utilized $4$ layers in the decoder and encoder in UNet and $6$ MoDL blocks. We used a batch size of $2$ during training for both the global and local cases. Furthermore, for the data-consistency term, we used a tolerance of $10^{-5}$ in CG and a $\mu/\nu$ ratio of 0.1. Also, we chose the regularization weight $\lambda$ as $10^{-9}$ for LONDN-MRI, unless specified otherwise.
@@
For the second MoDL architecture, we used the recent state-of-the-art denoising network DIDN [45, 46]. Due to the high complexity of the DIDN network, we first pre-trained it on the larger (global) dataset (learning rate, etc., similar to the UNet case) before adapting the weights within LONDN-MRI for each scan. This is an alternative to constructing shallower versions of a network for local adaptation.
The ADAM optimizer was utilized for training, with a learning rate of $5\times 10^{-5}$ in LONDN-MRI. We used $6$ iterations of MoDL with the DIDN denoiser for which we used $3$ down-up blocks (DUBs). The number of epochs for training was $30$ in LONDN-MRI. The remaining training parameters were chosen similarly as in the previous UNet-based case. Using a pre-trained state-of-the-art denoiser allows the local adaptation to converge faster.
@@ IV-B Results and Comparisons
When varying the size of the training set, the global approach was trained on the full set each time, whereas the local approach performed training on small subsets of $30$ training pairs selected from the larger datasets. The iterations of the LONDN-MRI scheme quickly improve reconstruction performance, and even with only 2 LONDN-MRI alternations, the PSNR values begin approaching the oracle setting.The LONDN schemes (oracle or iterative) consistently outperform the globally trained networks across the different training set sizes considered. We note that the results for the globally trained model with many ($6000$) training scans match closely the LONDN-MRI results, when LONDN-MRI uses a smaller overall training set ($3000$ scans) for neighbor search. This illustrates the potential of our approach with limited training data, when compared with models trained on larger sets.
Figure 2 compares the SSIM and HFEN reconstruction metrics using bar graphs, where a similar trend is observed as with PSNR.
Figs. 4 and 5 show images reconstructed by different methods at 8x and 4x undersampling, respectively.
The LONDN-MRI reconstructions (either iterative or oracle) show fewer artifacts, sharper features, and fewer errors than the global MoDL and initial aliased reconstructions. The iterative LONDN-MRI results are also quite close to the oracle result.
@@ IV-C Behavior of LONDN-MRI
$\text{NMA}:=\frac{100}{|\mathcal{T}|}\sum_{r\in\mathcal{T}}\frac{|\hat{C}_{r}\cap C_{r}^{*}|}{k},$ (8)
@@
Convergence of Loss in Bilevel Optimization: Next, we study the behavior of the alternating LONDN-MRI algorithm as a heuristic for the bilevel optimization formulation in (7). Here, we used an MoDL network with the UNet denoiser and $k=30$ training pairs were chosen (from 3000 cases) in the local dataset in each iteration of LONDN-MRI. The UNet weights were randomly initialized to begin with, and the neighbor search in the first iteration of LONDN-MRI was performed using $\bm{x}^{0}$ and correspondingly generated aliased training images. Fig. 11 plots the upper-level loss in (7) (in a root mean squared error form) after each iteration of LONDN-MRI for a test image. Here, we ran many iterations to verify convergence. We observe that the loss changes very little after a few iterations and stabilizes. This matches with the behavior of the neighbor search accuracy bar plots. The result indicates that the proposed alternating scheme could be a reasonable heuristic for reducing the loss in the challenging problem (7). Finally, we compare the loss values in Fig. 11 with an oracle loss, where the upper-level loss in (7) is computed using the ground truth test image and its $k$ nearest neighbors. The loss values in LONDN-MRI converge very close to the oracle loss, indicating potential for our scheme.
@@
Time Consumption Trade-offs: To further understand the time efficiency of our method across different neighborhood sizes for practical applicability, we conducted comparative analyses using three models: an image-domain UNet denoiser, MODL with UNet denoiser, and the MODL with DIDN denoiser. The experiments were run on an NVIDIA GeForce RTX A5000 GPU. The PSNR vs. runtime trade-offs depicted in Figure 13 shed light on the time consumption for each model configuration. It is observed that some decrease in the number of neighbors leads to reduced time consumption without significantly compromising image quality. In addition, the results show the effectiveness of starting with a pre-trained DIDN model to improve the reconstruction, as it enhances the efficiency of the reconstruction process, reducing it to order of seconds.